Week of 160711

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maria Alandes (chair and minutes), Renato Santana (LHCb), Cedric Serfon (ATLAS), Maarten Litmaath (ALICE), Andrea Manzi (MW Officer), Vincent Brillault (Security)
  • remote: Daniele Bonacorsi (CMS), Eric Lancon (BNL), Francesco Noferini (CNAF), Dave Mason (FNAL), Rolf Rumler (IN2P3), Dimitri (KIT), Christian (NDGF), Kyle Gross (OSG), Pepe Flix (PIC), Tiju Idiculla (RAL), Di Qing (TRIUMF)

Experiments round table:

  • ATLAS reports (raw view) -
    • Activities:
      • T0 spillover ran at BNL on the biggest run (38 hours)
    • Problems:
      • Problems affecting EOS :
        • Namespace server crashed, then bad space reported after
        • Bad throughput rate from EOS to T1_TAPE for some files due to switches
      • Some T1s getting full. Deletion campaign of derived data is being done + rebalancing of data

Maria asks whether there is any ticket tracking the EOS switch issues. Cedric replies that there is no ticket. This is being followed up internally with mails and meetings. It is suggested to ask the storage team to enter a report in the minutes twiki giving some details about the switch issues.

Maarten reports that GGUS:122743 (CERN-PROD nodes cannot access the KV cache) opened during the weekend should be also be monitored by WLCG Ops and recorded in the minutes twiki.

  • CMS reports (raw view) -
    • (Daniele at the CMS MB today - will try to connect, depending on where we are with the agenda)
    • Excellent LHC performance continues to keep the T0 busy for CMS
      • quite some debugging and optimization has been done at EOS, FTS, PhEDEx levels. Reference ticket is GGUS:122415
      • T0-outbound traffic more healthy, plus thanks to internal CMS re-prioritizations we are increasing our ability to send out of T0 the most urgent data tiers
      • we need to work on higher transfer rates though, we are confident we can achieve more (e.g. on per-file transfer rates) with an ad-hoc task force in synergy with EOS/FTS friends

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Activity
      • Monte Carlo simulation, data reconstruction/stripping and user jobs on the Grid
    • Site Issues
      • T0:
      • T1: NTR

Sites / Services round table:

  • ASGC: NA
  • BNL: NTR
  • CNAF: On Wednesday morning (13/07) we have planned an intervention on LHCb storage (scheduled as at risk). ATLAS reported a problem which may be related to network issues (GGUS:122737), under investigation.
  • FNAL: NTR
  • GridPP: NA
  • IN2P3: NTR
  • JINR: NA
  • KISTI: NA
  • KIT: S storage controller was replaced during Monday 11.07 with a short interruption for ALICE.
  • NDGF: NTR
  • NL-T1: NA
  • NRC-KI: NA
  • OSG: NTR
  • PIC: NTR
  • RAL: NTR
  • TRIUMF: Main site power transformation was problematic, as a prevention all services and equipment were shut down around 10:00 (UTC time) today, will bring all services back when the problem is solved.

  • CERN computing services: NA
  • CERN storage services:
    • FTS upgrade tomorrow morning: OTG:0031506.
      • Andrea asks FNAL whether the upgrade of their FTS server is scheduled soon. The upgrade is needed by CMS to avoid problems with transfers to dCache sites that last more than 1h. FNAL replies that this will be done this week, in any case, most dCache sites have increased their timeouts from 1h to 8h, so the upgrade is not urgent.
    • EOS performance problems. Thanx to CS some new switches are being deployed to (at least partially) bypass faulty (misbehaving) switches. In parallel we are also preparing to have more gFTP doors. We expect this to be visible tomorrow. On Wednesday we will touch base with ATLAS and CMS.
    • EOS stability. Meeting with ATLAS (this morning). Agreed on investigating the situation: for the moment mainly understanding recent crashing (both sw side and checking client version).
  • CERN databases: CMSONR node unavailability: The enclosure needs replacement which will be done during the "end of July" machine development. In the meantime, the database is running on the other nodes without problem (and a spare node is available in the case the situation could not be solved easily).
  • GGUS: NA
  • Monitoring:
    • Draft reports for the June availability sent around
  • MW Officer: NTR
  • Networks: NA
  • Security: Vulnerability in perfSONAR (http://www.perfsonar.net/#20160707-security)

AOB:

Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2016-07-12 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback